Bridgeport Corporation has fixed costs of $4,360,300. It has a unit selling price of $9.10, unit variable costs of $4.40, and a target net income of $1,510,000.

Compute the required sales in units to achieve its target net income.

enter a number of units for the Required sales
Answer #1

Contribution margin=Sales-Variable cost

=9.1-4.4=$4.7 per unit

Target Contribution margin=Fixed cost+Target profits

=4,360,300+1,510,000=$5870300

Hence target sales=$5870300/4.7

=1,249,000 units